# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ V.8 Real-Time Time-Dependent DFT
A more accurate theory for simulating excited states was developed by Runge and Gross Runge and Gross (1984), which is named time-dependent density functional theory (TDDFT). Real-time TDDFT (rt-TDDFT) can investigate electron dynamics, such as optical absorption spectra Ren, Kaxiras, and Meng (2010), stopping power Maliyov, Crocombette, and Bruneval (2018), photocatalysis Ren et al. (2024), and field-induced transitions. However, its computational intensity makes it challenging for large systems. Thus, employing numerical atomic orbital basis sets is advantageous, as they significantly reduce the computational burden while allowing for detailed analysis of excited-state phenomena and their intricate behaviors. Meng and Kaxiras (2008)
@@
The propagation of electrons obeys the time-dependent Kohn-Sham (TD-KS) equation:
$\mathrm{i}\frac{\partial}{\partial t}\psi_{j}(\mathbf{r},t)=\hat{H}_{\mathrm{KS}}[n(\mathbf{r},t)]\psi_{j}(\mathbf{r},t).$ (123)
@@
The adiabatic approximation is applied, so the Hamiltonian depends only on the instantaneous charge density. For numerical atomic orbitals, the TD-KS equation can be expressed in matrix form Lian et al. (2018):
$\mathrm{i}\frac{\partial\mathbf{c}}{\partial t}=S^{-1}H\mathbf{c}.$ (124)
@@
where $S$ represents the overlap matrix, and $\mathbf{c}$ is the column vector of coefficients for the local basis. For simplicity, the subscripts denoting the band index and $\mathbf{k}$-points have been omitted.
In ABACUS, we use the Crank-Nicolson method to approximate the propagator:
$\mathbf{c}(t+\mathrm{d}t)=\frac{S(t+\mathrm{d}t/2)-\mathrm{i}\hbar H(t+\mathrm{d}t/2)\mathrm{d}t/2}{S(t+\mathrm{d}t/2)+\mathrm{i}\hbar H(t+\mathrm{d}t/2)\mathrm{d}t/2}\mathbf{c}(t),$ (125)
where $S(t+\mathrm{d}t/2)$ and $H(t+\mathrm{d}t/2)$ are calculated using a linear approximation. Since $H(t+\mathrm{d}t/2)$ depends on $\mathbf{c}(t+\mathrm{d}t)$, a self-consistent procedure is needed to perform the propagation.
Ehrenfest dynamics is adopted for ion-electron coupled systems. The Verlet algorithm is employed to calculate the ionic velocities and positions at each time step. Fig. 29 shows a basic procedural flowchart of the rt-TDDFT algorithm in ABACUS.
Set $\rho(t_{2})$ as initial $\rho(t_{1})$, i.e., $\rho(t_{2})\leftarrow\rho(t_{1})$Construct the Hamiltonian at time $t_{2}$ using $\rho(t_{2})$Construct $H_{\mathbf{k}}(t_{2})$ and $S_{\mathbf{k}}(t_{2})$$\psi_{\mathbf{k}}(t_{2})\leftarrow\exp\left[\mathrm{i}(t_{1}-t_{2})S_{\mathbf{k}}^{-1}H_{\mathbf{k}}\right]\psi_{\mathbf{k}}(t_{1})$Construct $\rho^{\mathrm{new}}(t_{2})$ using $\psi_{\mathbf{k}}(t_{2})$$||\rho^{\mathrm{new}}-\rho||<\eta?$$\rho\leftarrow(1-w)\rho+w\rho^{\mathrm{new}}$ and other post-processingYesNo
Figure 29: Flowchart of the rt-TDDFT algorithm in ABACUS. It is worth noting that the self-consistent iteration of the electron density $\rho(t)$ ensures stability, with convergence criteria $||\rho^{\mathrm{new}}-\rho||<\eta$. The final stage involves post-processing, calculating quantities like total energy using Harris functional, and extracting physical observables.
The intensity of optical absorption spectra can be expressed as the imaginary part of the dielectric function:
$S(\omega)=\frac{2\omega}{3\pi}\mathrm{Im}\{\mathrm{tr}(\alpha_{\mu,\mu}(\omega))\},$ (126)
and the dielectric function can be calculated through the change in the dipole moment $P_{\mu}$ induced by the external laser field:
$\alpha_{\mu,\nu}(\omega)=\frac{\int P_{\mu}(t)\mathrm{e}^{\mathrm{i}\omega t}\,\mathrm{d}t}{\int E_{\nu}(t)\mathrm{e}^{\mathrm{i}\omega t}\,\mathrm{d}t}.$ (127)
Different electric fields in the time domain are provided in ABACUS, such as the Heaviside step function and Gaussian function. A Gaussian-type electric field is often used:
$E(t)=E_{0}\cos(2\pi ft+\phi)\exp(-\frac{(t-t_{0})^{2}}{2\sigma^{2}}).$ (128)
The Hamiltonian, including electric fields, has two forms: length gauge and velocity gauge. Within the length gauge, the extra Hamiltonian term is $V_{\mathrm{ext}}$:
$V_{\mathrm{ext}}=e\mathbf{E}(t)\cdot\mathbf{r}.$ (129)
@@
However, due to the periodic nature of the unit cell, this method will break the translational invariance of the electric potential. To address this, a sawtooth field in the spatial domain is used:
@@
where $\mu=x,y,z$, $L_{\mu}$ is the length of the unit cell along $\mu$, and $\epsilon\to 0$. To avoid divergence, the charge density must vanish in the region $-\epsilon<x_{\mu}<\epsilon$. Ideally, this region should be set as a vacuum layer. Therefore, the length gauge can only be used for finite systems theoretically.
Another method is the velocity gauge, which introduces a vector potential to simulate the laser field:
$\mathbf{A}=-c\int\mathbf{E}\,\mathrm{d}t.$ (131)
@@
$H_{\mathbf{k}}=\frac{1}{2m}\left(\hbar\mathbf{k}-\frac{e}{c}\mathbf{A}\right)^{2}.$ (132)
Below, we had a benchmark calculation for the optical absorption spectra of an ozone molecule ($\mathrm{O}_{3}$). The results are shown in Fig. 30.
Figure 30: Optical absorption spectra of the ozone molecule calculated using real time TDDFT in ABACUS. The results obtained using the length gauge and velocity gauge are consistent with each other and show reasonable agreement with experimental data Mason et al. (1996) and ab initio results from SIESTA Ren, Kaxiras, and Meng (2010).
Under normal temperature and pressure conditions, the electronic excitation energy is much smaller than the semiconductor band gap or chemical bond dissociation energy. The conventional Born-Oppenheimer molecular dynamics can be used in such cases. However, when electrons are excited by high-energy photons or extremely high temperatures, there is a strong coupling between electron motion and atomic nucleus motion, causing the Born-Oppenheimer approximation no longer valid. Typical applications include light-matter interactions and warm dense matter. TDDFT as an excitation state calculation method, can achieve high computational accuracy by choosing appropriate exchange-correlation potentials. It is widely used in areas such as light-controlled material phase transitions Liu et al. (2023b), time-dependent evolution of excited-state quasi-particles Sun et al. (2021), and photochemical reaction mechanism analysis You et al. (2024).
In photocatalysis, the timescale of electron motion typically ranges from femtoseconds (fs) to picoseconds (ps). For example, the lifetime of hot electrons captured by wet electron states on TiO2 surfaces, where water molecules are adsorbed, is usually less than 15 fs Onda et al. (2005). Correspondingly, the chemical reactions driven by photogenerated carriers also occur within the fs to ps timescale. Photocatalysis can also induce localized hotspot effects, significantly impacting product selectivity and efficiency. The interaction among photons, phonons, and molecules at single-atom sites is crucial for understanding energy conversion efficiency and molecular reaction selectivity in practical photocatalytic systems.
@@
Focusing on the photocatalysis, Liu et al. Ren et al. (2024) studied the dissociation process of water molecules under thermal and photoexcitation conditions using Ni single atoms supported on CeO2, based on the TDDFT functionality of ABACUS, as shown in Fig. 31. The ground-state electronic density shows that when water molecules adsorb on Ni single atoms, due to the presence of oxygen-rich defects, the anti-bonding state formed by the hybridization of the Ni atom d orbitals and H2O molecular orbitals is occupied (Fig. 31a), leading to weak adsorption strength of H2O and a correspondingly low dissociation barrier. Under thermal excitation at 600 K, H2O dissociates into H and OH, causing the oxygen vacancies near the Ni exposure site to be covered by OH, which in turn deactivates the catalyst. However, under photoexcitation, the Ni atom site accumulates holes transferred to the H2O molecule, weakening the adsorption between Ni and H2O. As a result, the dissociation of H2O to form O does not fill the oxygen vacancies around the Ni site, ensuring the stability of the catalyst. This mechanism has also been confirmed experimentally.
In another study, Liu et al. Chen et al. (2024b) compared the dissociation of CO2 driven directly by hot carriers generated via plasmonic excitation in metal clusters with the thermochemical reduction of CO2 by H species. They used TDDFT to elucidate the mechanisms and characteristic timescales of both processes and provided experimental evidence from in situ infrared spectroscopy. Additionally, TDDFT is suitable for analyzing charge transfer pathways, proton-coupled electron transfer (PCET), plasmonic processes, and other phenomena in photocatalytic reactions.
@@
Figure 31: TDDFT applications on photocatalysis Ren et al. (2024). (a) Projected density of states for H2O on Ni single atom site load on CeO2. (b) Bond length of thermal dissociation of H2O under 600 K. (c-d) Bond length of light-induced dissociation of H2O. (Adapted with permission from Nat. Commun. 15, 4675 (2024). Copyright 2024 Springer Nature.)
VI ABACUS in the OpenLAM Project
As the collection of quantum mechanical data progressively encompasses the entire periodic table, the Deep Potential team has launched an ambitious project named the Large Atomic Model (LAM) based on the practice of DPA pre-training model Zhang et al. (2024c, d). Given its openness and the crucial role of transparency in its progress, we call it the OpenLAM project ope (2024a).
While extensive ab initio data repositories are available, precise sampling for specific application scenarios remains vital for the pursuit of the model’s accuracy and transferability. To support the development of OpenLAM, the ABACUS team has launched the NOVA (Nourishing OpenLAM via ABACUS) project. This project is dedicated to supplying OpenLAM with affordable, high-precision first-principles data across diverse practical contexts. To this end, the ABACUS team help the users choose highly accurate pseudopotentials for various scenarios of OpenLAM through systematic precision testing, which will be introduced below in Sec. VI.1. By implementing heterogeneous acceleration (Sec. IV.2) and aggressively optimizing memory usage, along with economical hardware, the ABACUS team has significantly reduced the overall cost of data production. ABACUS has thus far generated a wealth of data in multiple application fields such as semiconductors, alloys, and Superhydrides to facilitate the ongoing refinement of OpenLAM within these sectors.
VI.1 APNS Project
Advanced computational solid-state datasets developed in recent decades—such as Materials ProjectJain et al. (2020), Materials CloudTalirz et al. (2020), and Computational Materials RepositoryLandis et al. (2012)—extensive literature has clarified correct DFT practices and explored ways to fine-tune the balance between precision and efficiency. In recent years, professional discussions on how to validate DFT code implementations are becoming more prevalentTalman (1978); Toyoda and Ozaki (2010); Wang et al. (2019b); Lejaeghere et al. (2016). A comprehensive and systematic verification of ABACUS remains essential. This measure is imperative to maintain efficient and reliable software development and, by extension, to guarantee optimal efficiency and precision in high-throughput computations.
Technically, contemporary DFT codes for periodic system calculations have greatly benefited from the pseudopotential formulation and the collaborative nature of open-source pseudopotential development. Numerous open-source codes for generating pseudopotentials have been released, enabling users to tailor pseudopotentials to their specific requirements Schlipf and Gygi (2015); Holzwarth et al. (2022). This customization also extends to the numerical atomic orbitals used in ABACUS for LCAO calculations. However, in practical DFT calculations, basis sets are ubiquitously employed to approximate physically acceptable functions—ranging from wavefunctions and charge densities to any function transitioning between real space and its Fourier-transformed counterpart. Thus, it is crucial to validate the size of basis as well as the quality of pseudopotentials and orbitals for particular systems before practical application, as those latter encapsulate atomic characteristics that indirectly influence critical physical properties of interest, such as elastic modulus and band gap, while the former controls the level describing atomic characters. We initiated the “ABACUS Pseudopotential Numerical Atomic Orbital Square” (APNS) project. This suite of workflows facilitates automated, high-throughput, and even user-interactive extended testings. It encompasses tests for pseudopotentials and numerical atomic orbitals across ideal, practical, or user-defined systems, accommodating both PW and LCAO calculations. Additionally, APNS is designed to offer a database where users can review test outcomes and download the appropriate pseudopotential and orbital files.
The systematic test workflow is depicted in Fig. 32. An additional front-end interface of this workflow is currently in development, the established back-end workflow has been employed to conduct efficiency and precision tests on a variety of norm-conserving pseudopotential sets. These pseudopotential sets include SG15SG15 ONCV Potentials , PseudoDojoPseudo Dojo , PD03PWmat-website , PD04PWmat-website (along with their versions and semi-core configurations, where available), as well as certain Goedecker-Teter-Hutter pseudopotentialsKrack (2005) utilized in CP2K, and three widely-used types of ultrasoft pseudopotentials (pslibrary 0.3.1 and 1.0.0theos-pp , GBRV v1.5Garrity et al. (2014); GBRV pseudopotentials ). This workflow has also played a role in developing various ABACUS features, such as DFT+U and support for ultrasoft pseudopotentials.
@@ VI.1 APNS Project
Figure 32: Workflow-chart of ABACUS Pseudopotential-Numerical atomic orbital Square automatic, high-throughput workflow incorporating structure generation, download and management (interfaced with online open crystal structure databases Materials ProjectJain et al. (2020), OptimadeAndersen et al. (2021), Crystallography Open Database (COD)Gražulis et al. (2009), etc.), symmetry analysis (interfaced with SeeK-pathTogo, Shinohara, and Tanaka (2024); Hinuma et al. (2017)), modular and extensible software input generation (ABACUS and Quantum ESPRESSOGiannozzi et al. (2020)) and remote high-throughput computing support.
As mentioned, the kinetic energy cutoff in plane-wave (PW) calculations is a major determinant of result precision, yet excessively high cutoff values are impractical for routine calculations. Therefore, tests were designed to find the best balance between efficiency and accuracy for each pseudopotential construction, family, and version. Technically, this involves conducting convergence tests on various parameters and performing Equation-of-States (EOS) tests using the converged cutoff values.
The kinetic cutoff energy of plane wave basis convergence tests are done on the most stable crystal phases collected from Materials ProjectTalirz et al. (2020), Crystallography Open Database (COD)Gražulis et al. (2009), OptimadeAndersen et al. (2021), etc. Keeping all other parameters unchanged, the value of cutoff varies from 20 to 300 Ry. For various kinds of pseudopotentials, the convergence of the PW kinetic energy cutoff is tested on three properties: Kohn-Sham energy, lattice pressure (trace of stress tensor) and a simplified version of band structure similarityPrandini et al. (2018)
$\eta_{v}(A,B)=\sqrt{\frac{\sum_{n\mathbf{k}}{\tilde{f}_{n\mathbf{k}}(e_{n\mathbf{k}}^{A}-e_{n\mathbf{k}}^{B})^{2}}}{\sum_{n\mathbf{k}}{\tilde{f}_{n\mathbf{k}}}}},$ (133)
where $e^{A}_{n\mathbf{k}}$ is the energy of the $n$-th band at kpoint $\mathbf{k}$ of calculation with DFT setup $A$ (so as for $B$), $\tilde{f}_{n\mathbf{k}}=\sqrt{f_{n\mathbf{k}}^{A}f_{n\mathbf{k}}^{B}}$, is the geometrical average occupation between calculation with two cutoff values.
@@
Figure 33: ABACUS PW calculation efficiency and precision tests on SG15 1.0 pseudopotentials. Number below the element symbol labels the converged kinetic energy cutoff ($E^{\mathrm{kin}}_{\mathrm{c}}$) under three thresholds: Kohn-Sham energy (< 1 meV/atom), pressure (< 0.1 kbar) and band similarity (< 10 meV). Large $E^{\mathrm{kin}}_{\mathrm{c}}$ which are enclosed by brackets are result from the oscillation or slow convergence of pressure. Color of each block indicates the $\Delta$ value calculated with converged $E^{\mathrm{kin}}_{\mathrm{c}}$ averaged over Bravais lattices BCC, FCC and Diamond. Blocks of elements whose pseudopotentials are not available are left empty.
@@
EOS tests are done not-spin-polarized, with k-point being sampled with spacing 0.06 Bohr-1 in Brillouin zone for all elements. Volumes of crystal vary from 94% to 106% with stepsize 2%. The kinetic energy cutoff of the PW basis is taken as the maximum among converged values of all elements involved in the system tested. $\Delta$ valueLejaeghere et al. (2014), which is defined as,
@@
is used as a scalar to indicate the difference between the EOS profile calculated by pseudopotential calculation and the all-electron one. The continuous integral in Eq. (134) is done on the Birch-Murnaghan equationBirch (1947)

### v3

@@ IV.7 Real-Time Time-Dependent DFT
DFT is effective in predicting ground-state properties but faces challenges when dealing with excited states. A more accurate theoretical framework for simulating excited states was proposed by Runge and Gross, Runge and Gross (1984) known as time-dependent density functional theory (TDDFT). Real-time TDDFT (rt-TDDFT) enables the investigation of electron dynamics, including applications in optical absorption spectra, Ren, Kaxiras, and Meng (2010) stopping power, Maliyov, Crocombette, and Bruneval (2018) photocatalysis, Ren et al. (2024) and field-induced transitions. However, its computational cost presents a significant hurdle for large-scale systems. In this context, the use of numerical atomic orbital (NAO) basis sets helps reduce computational expenses while still permitting detailed analysis of excited-state phenomena and their complex behaviors. Meng and Kaxiras (2008)
@@
The propagation of electrons obeys the time-dependent Kohn-Sham (TD-KS) equation
$\mathrm{i}\frac{\partial}{\partial t}\psi_{j}(\mathbf{r},t)=H\psi_{j}(\mathbf{r},t).$ (106)
@@
Here the Hamiltonian $H$ depends only on the electron density at time $t$ and we adopt the adiabatic approximation. For numerical atomic orbitals, the TD-KS equation can be expressed in the matrix form Lian et al. (2018)
$\mathrm{i}\frac{\partial\bm{c}}{\partial t}=\bm{S}^{-1}\bm{H}\bm{c}.$ (107)
@@
where $\bm{S}$ represents the overlap matrix, and $\bm{c}$ is the column vector of coefficients for the local basis. For simplicity, the subscripts denoting the band index and $k$-points have been omitted. We adopt the first-order Crank-Nicolson method to approximate the propagator
$\bm{C}(t+\mathrm{\Delta}t)=\frac{\bm{S}(t+\mathrm{\Delta}t/2)-\mathrm{i}\bm{H}(t+\mathrm{\Delta}t/2)\mathrm{\Delta}t/2}{\bm{S}(t+\mathrm{\Delta}t/2)+\mathrm{i}\bm{H}(t+\mathrm{\Delta}t/2)\mathrm{\Delta}t/2}\bm{C}(t),$ (108)
where $S(t+\mathrm{\Delta}t/2)$ and $H(t+\mathrm{\Delta}t/2)$ are computed using a linear approximation. Since $H(t+\mathrm{\Delta}t/2)$ depends on $\mathbf{C}(t+\mathrm{\Delta}t)$, a self-consistent procedure is needed to perform the propagation. The approach to approximating the propagator ensures its time-reversal invariance, conserves the total energy effectively, and preserves the orthogonality of the propagating wave functions. In addition, the Ehrenfest dynamics is adopted for ion-electron coupled systems and the Verlet algorithm is employed to calculate the ionic velocities and positions at each time step.
The time-dependent electric field $\mathbf{E}(t)$ is introduced into the Hamiltonian to simulate the interactions between laser field and materials, but there are different ways to implement the electric field. First, within the length gauge, the electric field potential term in Hamiltonian takes the form of
$V_{\mathrm{efield}}(\mathbf{r},t)=\mathbf{E}(t)\cdot\mathbf{r}.$ (109)
@@
However, the periodic nature of the cell breaks the translational invariance of the electric potential. To address this, a sawtooth field in the spatial domain takes the form of
@@
where $\mu=x,y,z$, $L_{\mu}$ is the length of cell along $\mu$, and $\epsilon\to 0$. To avoid divergence, the charge density must vanish in the region $\epsilon<x_{\mu}<\epsilon$. Ideally, this region should be set as a vacuum layer. Therefore, the length gauge can only be used for finite systems. The second method is the velocity gauge, which introduces a vector potential $\mathbf{A}(t)$ to simulate the laser field as
$\mathbf{A}(t)=-\int\mathbf{E}(t)\,\mathrm{d}t.$ (111)
@@
$H_{\mathrm{kinetic}}=\frac{1}{2}\Big{[}-\mathrm{i}\nabla+\mathbf{A}(t)\Big{]}^{2}.$ (112)
This method maintains the periodicity of the system. Third, while rt-TDDFT in periodic systems often relies on the velocity gauge to maintain periodicity under external fields, NAO implementations face a critical limitation: the position-dependent phase variations induced by the vector potential are neglected, leading to errors especially for properties like current. To overcome this challenge, Zhao et al. Zhao and He (2025) introduces a hybrid gauge in ABACUS that incorporates both the electric field and the vector potential, explicitly preserving phase information within atomic orbitals. Benchmark results confirm that this approach fully rectifies the inaccuracies of the pure velocity gauge in NAO-based rt-TDDFT, delivering reliable and precise simulations of ultrafast dynamics.
@@
Based on the rt-TDDFT functionality of ABACUS, Liu et al. Ren et al. (2024) studied the dissociation process of water molecules under thermal and photoexcitation conditions using Ni single atoms supported on CeO2, as shown in Fig. 26. The ground-state electronic density shows that when water molecules adsorb on Ni single atoms, due to the presence of oxygen-rich defects, the anti-bonding state formed by the hybridization of the Ni atom $d$ orbitals and H2O molecular orbitals is occupied (Fig. 26a), leading to weak adsorption strength of H2O and a correspondingly low dissociation barrier. Under thermal excitation at 600 K, H2O dissociates into H and OH, causing the oxygen vacancies near the Ni exposure site to be covered by OH, which in turn deactivates the catalyst. However, under photoexcitation, the Ni atom site accumulates holes transferred to the H2O molecule, weakening the adsorption between Ni and H2O. As a result, the dissociation of H2O to form O does not fill the oxygen vacancies around the Ni site, ensuring the stability of the catalyst. This mechanism has also been confirmed experimentally. In another study, Liu et al. Chen et al. (2024b) compared the dissociation of CO2 driven directly by hot carriers generated via plasmonic excitation in metal clusters with the thermochemical reduction of CO2 by H species. They used TDDFT to elucidate the mechanisms and characteristic timescales of both processes and provided experimental evidence from in situ infrared spectroscopy.
@@
Figure 26: Real-time TDDFT applications on photocatalysis Ren et al. (2024). (a) Projected density of states for H2O on Ni single atom site load on CeO2. (b) Bond length of thermal dissociation of H2O under 600 K. (c-d) Bond length of light-induced dissociation of H2O. (Adapted with permission from Nat. Commun. 15, 4675 (2024). Copyright 2024 Springer Nature.)
V ABACUS in the OpenLAM Project
As the collection of quantum mechanical data progressively encompasses the entire periodic table, the Deep Potential team has launched an ambitious project named the Large Atomic Model (LAM) based on the practice of DPA pre-training model Zhang et al. (2024c, d). Emphasizing open scientific collaboration, this initiative has been formally designated as the OpenLAM project ope (2024a), reflecting its commitment to transparent development and community-driven progress. ABACUS serves as the cornerstone computational platform for generating cost-effective yet high-fidelity first-principles data to support OpenLAM’s development. Through heterogeneous algorithm optimization specifically tailored for plane-wave basis solutions of the Kohn-Sham equations, the software achieves remarkable efficiency gains in data generation.
V.1 APNS Project
For those DFT codes that require pseudopotentials, several open-source codes for generating pseudopotentials are available, enabling users to tailor pseudopotentials to their specific requirements Schlipf and Gygi (2015); Holzwarth et al. (2022). For LCAO calculations using ABACUS, it is crucial to validate both the NAO basis set and the pseudopotentials before practical applications. To address the above issues, we have initiated the "ABACUS Pseudopotential Numerical Atomic Orbital Square" (APNS) project.
Systematic verification of ABACUS remains essential in the APNS project to ensure reliable software development and maintain efficient high-throughput calculations. This aligns with growing research on validating DFT codes Talman (1978); Toyoda and Ozaki (2010); Wang et al. (2019b); Lejaeghere et al. (2016). Major databases like Materials Project Jain et al. (2020), Materials Cloud Talirz et al. (2020), and Computational Materials Repository Landis et al. (2012) have improved DFT practices. These datasets help balance accuracy and efficiency in DFT calculations while establishing code implementation standards. The APNS workflow enables automated high-throughput testing with optional user interaction. It evaluates pseudopotentials and NAO basis set across predefined and customizable systems, compatible with both PW and LCAO methods. The system also maintains a results database for test verification and validated pseudopotential/orbital file downloads. Here we list some examples done in the APNS project.
Fig. 27 demonstrates the standardized testing workflow, where the workflow has conducted efficiency and accuracy evaluations for multiple norm-conserving pseudopotential families including SG15 SG15 ONCV Potentials , PseudoDojo Pseudo Dojo , PD03/PD04 PWmat-website (with version variations and semi-core configurations), CP2K’s Goedecker-Teter-Hutter pseudopotentials Krack (2005), and three ultrasoft types (pslibrary 0.3.1/1.0.0 theos-pp , GBRV v1.5 Garrity et al. (2014); GBRV pseudopotentials ). This workflow has additionally supported ABACUS feature development such as DFT+U implementation and ultrasoft pseudopotential compatibility.
@@ V.1 APNS Project
Figure 27: Workflow-chart of ABACUS Pseudopotential-Numerical atomic orbital Square automatic (APNS), high-throughput workflow incorporating structure generation, download and management (interfaced with online open crystal structure databases Materials Project Jain et al. (2020), Optimade Andersen et al. (2021), Crystallography Open Database (COD) Gražulis et al. (2009), etc.), symmetry analysis (interfaced with SeeK-path Togo, Shinohara, and Tanaka (2024); Hinuma et al. (2017)), modular and extensible software input generation (ABACUS and Quantum ESPRESSOGiannozzi et al. (2020)) and remote high-throughput computing support.
The PW kinetic energy cutoff critically determines calculation accuracy, but excessively high values hinder computational efficiency. To address this trade-off, we established a systematic protocol for optimizing pseudopotentials through convergence testing across key parameters. This process involves two-phase validation. First, convergence testing of cutoff energies (20-300 Ry range) using stable crystal structures from Materials Project Talirz et al. (2020), Crystallography Open Database Gražulis et al. (2009), and Optimade Andersen et al. (2021). Second, equation of state (EOS) verification with optimized cutoffs. Three physical properties, including Kohn-Sham energy, lattice pressure (stress tensor trace), and band structure similarity metrics Prandini et al. (2018) were used as convergence criteria across pseudopotential types.
@@
Figure 28: Efficiency and precision tests on SG15 1.0 pseudopotentials using PW basis of ABACUS. Number below the element symbol labels the converged kinetic energy cutoff ($E^{\mathrm{kin}}_{\mathrm{c}}$) under three thresholds: Kohn-Sham energy (< 1 meV/atom), pressure (< 0.1 kbar) and band similarity (< 10 meV). Large $E^{\mathrm{kin}}_{\mathrm{c}}$ values which enclosed by brackets are results from the oscillation or slow convergence with respect to pressure. Color of each block indicates the $\Delta$ value calculated with converged $E^{\mathrm{kin}}_{\mathrm{c}}$ averaged over BCC, FCC and Diamond structures. Blocks of elements whose pseudopotentials are not available are left empty.
@@
EOS tests were performed using non-spin-polarized calculations, with $k$-sampled at a spacing of 0.06 Bohr-1 in the Brillouin zone for all elements. The crystal volumes $V$ range from 94% ($V_{m}$) to 106% ($V_{M}$) with a 2% step size. The kinetic energy cutoff for the plane-wave (PW) basis is set to the maximum of the converged values for all elements in the tested system. The $\Delta$ value Lejaeghere et al. (2014), defined as
@@
serves as a scalar indicator to quantify the difference between EOS profiles derived from pseudopotential and all-electron calculations (denoted by $a$ and $b$, respectively). The continuous integral in Eq. (113) is evaluated over the curve fitted using the Birch-Murnaghan equation Birch (1947), where $E_{a}(V)$ and $E_{b}(V)$ represent the Kohn-Sham energies relative to their equilibrium values.
We constructed three Bravais lattices for each element, including BCC (Body-Centered-Cubic), FCC (Face-Centered-Cubic), and diamond structures. We calculate $E$-$V$ datasets and compare them with the all-electron results reported in Ref. Wang et al., 2019b. The converged kinetic energy cutoffs and corresponding $\Delta$ values for all available elements in the SG15 1.0 pseudopotential Schlipf and Gygi (2015) are shown in Fig. 28.
For the NAO basis set, we conducted precision tests. Similar to plane-wave (PW) calculations, precision is indicated by the EOS; additionally, the energy difference at the minimum of the EOS profile can serve as an indicator of basis set completeness. For ABACUS LCAO calculations, reliability verification tests for any given pseudopotential are often hindered by the relatively complex and time-consuming NAO generation procedure, which requires sophisticated experience and insights. This makes a rapid, highly automated orbital generation code necessary. Furthermore, as the precision and efficiency of NAOs are expected to improve continuously, a platform for orbital generation algorithm development is needed. Therefore, an open-source subproject, “ABACUS-ORBGEN”, has been included as a development component of APNS. The current implementation remains based on the algorithms proposed in Refs. Chen, Guo, and He, 2010a; Lin, Ren, and He, 2021b, but includes features such as automatic setting of orbital configurations for pseudopotentials generated with ONCVPSP codes, automated improvement of orbital transferability through a series of bond length scans and potential curve fitting tasks, and a quick orbital quality estimation function.
V.2 UniPero Model